OpenStudy (anonymous):

Find the inverse of F(x) =x-4. I really need someone to walk me through this.

jimthompson5910 (jim_thompson5910):

replace f(x) with y y = x - 4

jimthompson5910 (jim_thompson5910):

Swap x and y x = y - 4

jimthompson5910 (jim_thompson5910):

Then solve for y (ie isolate y) x = y - 4 x+4 = y - 4 + 4 x+4 = y y = x + 4

jimthompson5910 (jim_thompson5910):

So the inverse is \[\Large f^{-1}(x) = x+4\]

jimthompson5910 (jim_thompson5910):

So again, the basic outline is step 1) replace f(x) with y step 2) swap x and y step 3) solve for y to get the inverse

OpenStudy (anonymous):

Thank You so much my teacher really hasn't broken down what we do in class and we're having problems with our school curriculum so thank you I really appreciate this.:)

jimthompson5910 (jim_thompson5910):

you're welcome